Find the volume in liters of 40 kg of carbon tetrachloride, whose density is 1.60g/cm^3
1
Expert's answer
2015-06-30T06:50:48-0400
V (volume)=m (mass) /ρ(density)=(40*10^3)/1.60=25 *10^3 cm^3= 25 L